Determine the location and values of the absolute maximum and absoluteย 

minimum for the given function:ย 

๐‘“(๐‘ฅ) = (โˆ’๐‘ฅ + 2)

เฌธ

, ๐‘คโ„Ž๐‘’๐‘Ÿ๐‘’ 0 โ‰ค ๐‘ฅ โ‰ค 3


1
Expert's answer
2021-06-15T09:31:22-0400
"F(x)=(-x+2)^3"

"0\leq x\leq 3"

Find the first derivative


"F'(x)=((-x+2)^3)'=-3(-x+2)^2"

Find the critical number(s)


"F'(x)=0=>-3(-x+2)^2=0"

"x=2"

Critical number: "2"

"F(0)=(-2+0)^3=-8"




"F(3)=(-2+3)^3=1"

"F(2)=(-2+2)^3=0"

The function "F(x)" has the absolute maximum with value of "1" on "[0, 3]" at "x=3: Point(3,1)."


The function "F(x)" has the absolute minimum with value of "-8" on "[0, 3]" at "x=0: Point(0,-8)."